OpenStudy (anonymous):

you flip 3 coins what is the probability that you will get exactly 3 heads ?

OpenStudy (anonymous):

2*2*2 = 6 1/6

OpenStudy (anonymous):

1/2 *1/2 *1/2=1/8

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yeap the second one is right